The time it takes for an airplane production company to produce x number of their planes is given by 150x^-0.322(in thousand work-hours). How many work-hours will it take the production company to build their fifty-fifth airplane?
I do not understand this at all. Can someone please explain this to me?

Answer by lwsshak3(11628)

The time it takes for an airplane production company to produce x number of their planes is given by 150x^-0.322(in thousand work-hours). How many work-hours will it take the production company to build their fifty-fifth airplane?
***
just substitute for x the given number (55) airplanes then calculate the expression to get the answer in thousands of work hours.
150x^-0.322=150*55^-0.322)=150/(55^0.322)=41.28 thousand work hours